OpenStudy (idku):

INTEGRAL. I am practicing by making up problems. Don't always solve them, but in most the cases i do.

OpenStudy (idku):

\[\int\limits_{ }^{ }\sin^{-1}\sqrt{x}~dx\]

OpenStudy (idku):

So first I differentiate the inverse sine.

OpenStudy (idku):

the function,\[\sin(y)=\sqrt{x}\] the derivative, \[y'\times \cos(y)=\frac{1}{2\sqrt{x}}\]

OpenStudy (idku):

\[y'=\frac{1}{2\sqrt{x}\cos(y)}\]

OpenStudy (idku):

then if sin(y)=x/1, then the adjacent side is sqrt{1-x^2} So, \[y'=\frac{1}{2\sqrt{x-x^3}}\]

OpenStudy (idku):

So now, u=sin^-1(sqrt x) dv=1

OpenStudy (idku):

\[\int\limits_{ }^{ }\sin^{-1}(\sqrt{x})~dx=x\sin^{-1}x - \int\limits_{ }^{ }\frac{x}{2\sqrt{x-x^3}}~dx\]

OpenStudy (idku):

\[\int\limits_{ }^{ }\sin^{-1}(\sqrt{x})~dx=x\sin^{-1}x - \int\limits_{ }^{ }\frac{\sqrt{x}}{2\sqrt{1-x^2}}~dx\]

OpenStudy (idku):

@jim_thompson5910

OpenStudy (idku):

i was wondering if trig sub works

OpenStudy (idku):

oh if I sub x=cos(theta) i get it simplfied (or maybe that is even worse)

jimthompson5910 (jim_thompson5910):

I'm stuck honestly, but I'm still thinking it over

OpenStudy (idku):

I mean sin theta

OpenStudy (idku):

\[\int\limits\limits_{ }^{ }\frac{\sqrt{x}}{2\sqrt{1-x^2}}~dx\]\[x=\sin\theta\]\[dx=\cos\theta~d\theta\]

OpenStudy (idku):

\[\int\limits\limits_{ }^{ }\frac{\sqrt{\sin\theta}}{2\sqrt{1-\sin^2\theta}}\times \cos\theta~d\theta\]

OpenStudy (idku):

\[(1/2)\int\limits\limits_{ }^{ }\sqrt{\sin\theta}~d\theta\]

OpenStudy (idku):

but is this simpler to integrate?

jimthompson5910 (jim_thompson5910):

wolfram is getting this http://www.wolframalpha.com/input/?i=integral(arcsin(sqrt(x)))&t=crmtb01 but honestly I'm not sure how

jimthompson5910 (jim_thompson5910):

maybe it's using u-sub to say u = sqrt(x), du = dx/(2*sqrt(x)) ---> du = dx/(2u) ---> du = 2u*du but I'm not sure how to run with that

OpenStudy (idku):

from where i am now, it is deadly. by parts won't help, will just add more theta's to the trig functions. i did an incorrect trig sub i guess.

OpenStudy (idku):

wait, what if I said x=sin^2theta from there?

OpenStudy (idku):

i mean intead of just sin theta

jimthompson5910 (jim_thompson5910):

oops I meant to say dx = 2u*du but still not sure what to do with all that

OpenStudy (idku):

i think i might know it...

OpenStudy (idku):

\[\int\limits_{ }^{ }\frac{\sqrt{x}}{2\sqrt{1-x^2}}~dx\]\[u=\sqrt{1-x^2}\] sqrt{1-x^2}=u 1-x^2=u^2 -x^2=(u^2-1) no i fail to solve for sqrt(x) in terms of u. I get i.

OpenStudy (idku):

I will retry, maybe i can substitute u=sqrt{x} from the very beginning. tnx for trying

OpenStudy (idku):

\[\int\limits_{ }^{ }\frac{x}{2\sqrt{x}\sqrt{1-x^2}}~dx\]\[u=1-x^2\]\[x=\sqrt[{\LARGE 4}]{1-u}\]\[-\frac{1}{2}du=x~dx\]\[-\int\limits_{ }^{ }\frac{1}{\sqrt[4]{1-u}\sqrt{u}}~du\]\[-\int\limits_{ }^{ }\frac{1}{\sqrt[4]{u^2-u^3}}~du\]

OpenStudy (anonymous):

\[\int\arcsin(\sqrt x)\,dx\] Substituting \(u=\sqrt x\) gives \(u^2=x\), and \(dx=2u\,du\). \[\int\arcsin(\sqrt x)\,dx\stackrel{x\to u^2}{=}2\int u\arcsin u\,du\] Integrating by parts, you can set \[\begin{matrix}f=\arcsin u&&&dg=u\,du\\ df=\frac{du}{\sqrt{1-u^2}}&&&g=\frac{1}{2}u^2\end{matrix}\] \[\int u\arcsin u\,du=\frac{1}{2}u^2\arcsin u-\frac{1}{2}\color{red}{\int\frac{u^2}{\sqrt{1-u^2}}\,du}\] For the remaining integral, you can make the substitution \(u=\sin t\), then \(du=\cos t\,dt\). \[\color{red}{\int\frac{u^2}{\sqrt{1-u^2}}\,du}\stackrel{u\to\sin t}{=}\int\frac{\sin^2t\cos t}{\sqrt{1-\sin^2t}}\,dt=\int\sin^2t\,dt=\frac{1}{2}\int(1-\cos2t)\,dt\]

OpenStudy (anonymous):

@idku it looks like you tried IBP from the get-go. Setting that up like so: \[\begin{matrix} f=\arcsin(\sqrt x)&&&dg=dx\\ df=\frac{dx}{2\sqrt{x-x^2}}&&&g=x \end{matrix}\] and so \[\int \arcsin(\sqrt x)\,dx=x\arcsin(\sqrt x)-\frac{1}{4}\color{red}{\int\frac{x}{\sqrt{x-x^2}}\,dx}\] For this integral, complete the square in the radical and make a trigo sub. \[x-x^2=\frac{1}{4}-\left(x-\frac{1}{2}\right)^2\] Let \(x-\dfrac{1}{2}=\dfrac{1}{2}\sin p\), giving \(dx=\dfrac{1}{2}\cos p\,dp\). \[\begin{align*}\color{red}{\int\frac{x}{\sqrt{x-x^2}}\,dx}&\stackrel{x-\frac{1}{2}\to\frac{1}{2}\sin p}{=}\frac{1}{2}\int\frac{\left(\dfrac{1}{2}\sin p+\dfrac{1}{2}\right)\cos p}{\sqrt{\dfrac{1}{4}-\dfrac{1}{4}\sin^2p}}\,dp\\\\ &=\frac{1}{2}\int\left(\sin p+1\right)\,dp \end{align*}\]